OpenStudy (anonymous):

A new automobile is purchased for $20,000. If V = 20,000 (0.8)^x, gives the car's value after x years, approximately how long will it take for the care to be worth half of its purchase price? a. 3 years b. 4 years c. 5 years d. 6 years

OpenStudy (anonymous):

That's not an answer though.

OpenStudy (anonymous):

i no let me redo my work =(

OpenStudy (anonymous):

ha ha. I had the right work i just punched wrong button in calculator. V=20,000*(0.8)^x 10,000=20,000*(0.8)^x log(10,000/20,000)=x*log(0.8) x=[log(10,000/20,000)/log(0.8)] x=3.10628 =)